Please help to solve this problem. Thank you.
The volume of a cylinder is 175π. If the radius of the base of the cylinder is 5 in., what is the height of the cylinder.

Answer by scott8148(6628)

the volume of a cylinder is the product of the area of the base and the height

((pi)r^2)(h)=v __ ((pi)5^2)(h)=175(pi) __ (25pi)(h)=175pi __ 25h=175